please help me solve 4x+3y=13 if x is -2

Answer by ewatrrr(24785)


Hi
solving 4x+3y=13 for y when x = -2
The idea is to substitute that value for x and isolate y on one side of the equation
4x + 3y = 13
4*(-2) + 3y = 13 |substituting -2 for x
-8 + 3y = 13 | completing the multiplication 4*(-2)
3y = 21 |adding 8 to both sides of the equation
y = 7 | dividing both sides of the equation by 3
CHECKING our Answer
- 8 + 21 = 13